OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation (x+5)^2/81 - (y-3)^2/144 = 1.? @wio sorry last question!

OpenStudy (anonymous):

@watchmath

OpenStudy (watchmath):

\(a^2+b^2=81+144=225=15^2\) The vertex are at \((-14,0)\) and \((4,0)\). Hence the foci are at \((-14-15,0)=(-29,0)\) and \((4+15,0)=(19,0)\)

OpenStudy (watchmath):

sorry it is wrong. Give me a minute

OpenStudy (anonymous):

@watchmath that's okay! I understand the formula

OpenStudy (watchmath):

The center is at \((-5,3)\). So the foci are at (-20,3) and (10,3)

OpenStudy (anonymous):

so vertices at (4,3) and (-14,3)? @watchmath

OpenStudy (watchmath):

yes

OpenStudy (anonymous):

@watchmath THANK YOU SO MUCH! I think I get these problems now!(: